### v1

# Strict quantization for compact pseudo-Kähler manifolds and group actions

Abstract
The asymptotic results for Berezin-Toeplitz operators yield a strict quantization for the algebra of smooth functions on a given Hodge manifold. It seems natural to generalize this picture for quantizable pseudo-Kähler manifolds in presence of a group action. Thus, in this setting we introduce a Berezin transform which has a complete asymptotic expansion on the preimage of the zero set of the moment map. It leads in a natural way to prove that certain quantization maps are strict.
Keywords: CR manifolds, Toeplitz operators, group actions, star products
Mathematics Subject Classification: 32Vxx, 32A25, 53D50
1 Introduction
This paper is a continuation of [GH1], here we investigate further properties of the star product induced by Toeplitz operators, let us briefly recall some known results. In [BMS], the authors studied Berezin Toeplitz quantization for Hodge manifolds. In [MM2], see also references therein, Toeplitz operators for symplectic manifolds are introduced. In [GH1] we considered Berezin-Toeplitz quantization for a class of CR manifolds and investigated the properties of the corresponding star product, see also [GH2] for further results in this geometric setting. In [S2], additional properties for the Berezin Toeplitz star product are proved: it is shown that the properties of Berezin Toeplitz operators lead in a natural way to prove that certain quantization maps are strict. In this paper we aim to generalize these results to pseudo-Kähler manifolds in the presence of a group action.
Berezin-Toeplitz quantization has been studied extensively, giving rise to a vast literature. Without claiming to be exhaustive, let us mention a few references most relevant to the present discussion: [BoGu], [BPU], [C], [G], [HM], [KS], [P], [MM1], [MM2], [Z]. In particular, we refer to the surveys [S2], [W] for considerations about the induced star product. More recently, in [HHMS] and [Sh], the authors study Toeplitz operators associated with classical first-order pseudodifferential operators and apply them to problems in CR geometry. Eventually, in connection with the principle that “quantization commutes with reduction”, we refer to [Ma].
We now provide a more precise formulation of the concepts introduced above. Let $(X,\,T^{1,0}X)$ be a compact, connected and orientable CR manifold of dimension $12n+$ 1, $n\geq 1$, where $T^{1,0}X$ is a subbundle of the complexified tangent bundle $TX\otimes\mathbb{C}$. There is a unique sub-bundle $HX$ of $TX$ such that
$HX\otimes\mathbb{C}=T^{1,0}X\oplus T^{0,1}X$
which is called the horizontal bundle. Let $J\,:\,HX\rightarrow HX$ be the complex structure map given by
$J(u+\overline{u})=iu-i\overline{u}\,,$
for every $u\in T^{1,0}X$. By complex linear extension of $J$ to $TX\otimes\mathbb{C}$, the $i$-eigenspace of $J$ is $T^{1,0}X$.
Since $X$ is orientable, there always exists a real non-vanishing $1$-form
$\omega_{0}\in{C}^{\infty}(X,T^{*}X)$
so that $\langle\,\omega_{0}(x)\,,\,u\,\rangle=0$, for every $u\in H_{x}X$, for every $x\in X$. The one form $\omega_{0}$ is called contact form and it naturally defines a volume form $\mathrm{dV}_{X}$ on $X$. For each $x\in X$, we define a quadratic form on $HX$ with respect to $\omega_{0}$ by
${L}_{x}(U,V)=\frac{1}{2}\mathrm{d}\omega_{0}(JU,V),\qquad\forall\ U,V\in H_{x}X.$
Then, we extend ${L}$ to $HX\otimes\mathbb{C}$ by complex linear extension; for $U,V\in T^{1,0}_{x}X$,
${L}_{x}(U,\overline{V})=\frac{1}{2}\,\mathrm{d}\omega_{0}(JU,\overline{V})=-\frac{1}{2i}\,\mathrm{d}\omega_{0}(U,\overline{V}).$
The Hermitian quadratic form ${L}_{x}$ on $T^{1,0}_{x}X$ is called Levi form at $x$ (with respect to $\omega_{0}$). Hereafter we shall always assume that the Levi form is non-degenerate with $n_{-}$ negative and $n_{+}$ positive eigenvalues, $n_{+}+n_{-}=n$. The Reeb vector field $R\in{C}^{\infty}(X,TX)$ is defined to be the non-vanishing vector field determined by
$\omega_{0}(R)\equiv-1,\quad\mathrm{d}\omega_{0}(R,\cdot)\equiv 0\ \ \mbox{on $TX$}.$
There is a Reeb vector field such that the flow of it induces a transversal CR $\mathbb{R}$-action on $X$. Suppose that all orbits of the flow of $R$ are compact. Then $X$ admits a transversal CR circle action, which we denote by $e^{i\,\theta}\cdot$, where $\theta\in\mathbb{R}$.
Remark 1.1.
Note that the quotient, denoted by $M$, of $X$ by the circle action is a pseudo-Kähler manifold, see for example [BoGa]. Indeed, $M$ is endowed with an integral symplectic form $\omega$ such that
$2\,\pi^{*}\omega=\mathrm{d}\omega_{0}$
where $\pi\,:\,X\rightarrow M$ is the projection. The horizontal bundle $HX$ is the pullback of the tangent bundle $TM$ via $\pi$ and $J$ induces a complex structure $J_{M}$ on $M$. Thus,
$g:=\omega(\,\cdot,\,J_{M}\,\cdot\,)$
is easily seen to be a pseudo-Riemannian structure on $M$.
Vice versa, suppose that $(M,\omega,J_{M})$ is a compact pseudo-Kähler manifold with $[\omega]$ defining an integral cohomology class. Then, there exists a line bundle $(L,h)$ on $M$ such that the unique holomorphic connection compatible with $h$ has curvature $-i\,\omega$. Then, the circle bundle $X$ inside the dual of the line one is a contact manifold satisfying the assumption of this article.
To be consistent with [GH1] and [GH2] we prefer to work on $X$ and make explicit reference to $M$ later on.
Assume that $X$ admits an action of a compact connected Lie group $G$ of dimension $d$. In this work, we assume that the $G$ action preserves $\omega_{0}$ and $J$. That is, $g^{\ast}\omega_{0}=\omega_{0}$ on $X$ and $g_{\ast}J=Jg_{\ast}$ on $HX$, for every $g\in G$, where $g^{*}$ and $g_{*}$ denote the pull-back map and push-forward map of $G$, respectively.
Let $\mathfrak{g}$ denote the Lie algebra of $G$. For any $\xi\in\mathfrak{g}$, we write $\xi_{X}$ to denote the vector field on $X$ induced by $\xi$. That is,
$(\xi_{X}u)(x)=\frac{\partial}{\partial t}\left[u\left(e^{t\xi}\,x\right)\right]_{|_{t=0}},$
for any $u\in{C}^{\infty}(X)$.
[…]
Now, we define the quantizing space. Fix $g\in G$ and $x\in X$, for any $r\in\mathbb{N}$ let
$g^{*}:\Lambda^{r}_{x}(T^{*}X\otimes\mathbb{C})\rightarrow\Lambda^{r}_{g^{-1}\circ x}(T^{*}X\otimes\mathbb{C})$
be the pull-back map. Since $G$ preserves $J$, we have
$g^{*}:T^{*0,q}_{x}X\rightarrow T^{*0,q}_{g^{-1}\circ x}X\,$
for all $x\in X$, where the bundle $T^{*0,q}X$ is the $q$-th exterior power of $(T^{(0,1)}X)^{*}$, whose space of sections is denoted by $\Omega^{0,q}(X)$ and its elements are called $(0,q)$-forms. Thus, for $u\in\Omega^{0,q}(X)$, we have $g^{*}u\in\Omega^{0,q}(X)$. Put
$\Omega^{0,q}(X)^{G}:=\left\{u\in\Omega^{0,q}(X);\,g^{*}u=u,\ \ \forall g\in G\right\}\,.$
[…]
$L^{2}_{(0,q)}(X)^{G}:=\left\{u\in L^{2}_{(0,q)}(X):\,g^{*}u=u,\ \ \forall g\in G\right\}.$
Put
$({\rm Ker\,}\Box^{q}_{b})^{G}:={\rm Ker\,}\Box^{q}_{b}\cap L^{2}_{(0,q)}(X)^{G}.$
The $G$-invariant Szegő projection is the orthogonal projection
$S^{(q)}_{G}:L^{2}_{(0,q)}(X)\rightarrow({\rm Ker\,}\Box^{q}_{b})^{G}$
[…]
$S^{(q)}_{G}(x,y)\in{D}^{\prime}(X\times X,T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
be the distributional kernel of $S^{(q)}_{G}$, ${D}^{\prime}$ denotes the space of distribution sections, see Section 2.2 for notation.
In this paper, we will assume that $\Box^{q}_{b}$ has $L^{2}$ closed range. For simplicity we assume that $n_{+}\neq n_{-}$ and we fix $q=n_{-}$. Let $f\in C^{\infty}(X,\,\mathbb{C})$ be invariant under the circle action and the action of $G$, then the $G$-invariant Toeplitz operator is given by
$T^{G}[f]:=S^{(n_{-})}_{G}\circ f\cdot\circ S^{(n_{-})}_{G}:L^{2}_{(0,n_{-})}(X)\rightarrow({\rm Ker\,}\Box^{n_{-}}_{b})^{G}\,.$
The transversal circle action can be used to define the corresponding Fourier components. First, we assume that
$e^{i\theta}\cdot(g\,x)=g\,(e^{i\theta}\cdot x),\ \ \mbox{for every $x\in X$, $\theta\in[0,2\pi[$, $g\in G$}.$
Let $u\in\Omega^{0,q}(X)$ be arbitrary. Define
$Ru:=\frac{\partial}{\partial\theta}\bigr{(}(e^{i\theta})^{*}u\bigr{)}|_{\theta=0}\in\Omega^{0,q}(X).$
For every $k\in\mathbb{Z}$, let
$\Omega^{0,q}_{k}(X):=\left\{u\in\Omega^{0,q}(X);\,Ru=ik\,u\right\},\ \ q=0,1,2,\ldots,n,$
and
$\Omega^{0,q}_{k}(X)^{G}=\left\{u\in\Omega^{0,q}(X)^{G};\,Ru=ik\,u\right\},\ \ q=0,1,2,\ldots,n.$
[…]
$S^{(q)}_{G,k}:L^{2}_{(0,q)}(X)\rightarrow({\rm Ker\,}\Box^{q}_{b})^{G}_{k}$
[…]
$H^{G}(X):=({\rm Ker\,}\Box^{n_{-}}_{b})^{G}$
and
$H^{G}_{k}(X):=({\rm Ker\,}\Box^{n_{-}}_{b})^{G}_{k}\,.$
Eventually, for every $G$ and $S^{1}$ invariant function $f\in{C}^{\infty}(X)$, the $k$-th $G$-invariant Toeplitz operator is given by
$T^{G}_{k}[f]:=S^{(n_{-})}_{G,k}\circ f\cdot\circ S^{(n_{-})}_{G,k}:L^{2}_{(0,n_{-})}(X)\rightarrow H^{G}_{k}(X)\,.$ (1)
Definition 1.1.
The contact moment map associated to the form $\omega_{0}$ is the map $\mu:X\to\mathfrak{g}^{*}$ such that, for all $x\in X$ and $\xi\in\mathfrak{g}$, we have
$\langle\mu(x),\xi\rangle=\omega_{0}(\xi_{X}(x))\,.$ (2)
Let $b$ be the nondegenerate bilinear form on $HX$ such that
$b(\cdot,\cdot)=\mathrm{d}\omega_{0}(\cdot,J\cdot)\,.$ (3)
Let us denote $\mu^{-1}(0)$ by $Y$. Assume that 00 is a regular value of $\mu$, assume that the action of $G$ on $\mu^{-1}(0)$ is free and
[…]
where
$\underline{\mathfrak{g}}={\rm Span\,}(\xi_{X};\,\xi\in\mathfrak{g})$
and
$\underline{\mathfrak{g}}^{\perp_{b}}=\left\{v\in HX;\,b(\xi_{X},v)=0,\ \ \forall\xi_{X}\in\underline{\mathfrak{g}}\right\}\,.$
Then, $\mu^{-1}(0)$ is a $d$-codimensional submanifold of $X$. Let $Y:=\mu^{-1}(0)$ and let $HY:=HX\cap TY$. Note that if the Levi form is positive at $Y$, then (4) holds.
Remark 1.2.
Under notation as in Remark 1.1, the action of $G$ on $X$ defined above descends to a Hamiltonian and holomorphic action on $(M,\,\omega,\,J_{M})$. For each $\xi\in\mathfrak{g}$, we denote by $\xi_{M}$ the corresponding infinitesimal vector field on $M$. The moment map will be denoted by
$\Phi\,:\,M\rightarrow\mathfrak{g}^{*}\,.$
Vice-versa, we recall that given a pseudo-Kähler manifold $(M,\,\omega,\,J)$, whose $\omega$ defines an integral cohomology class. Here, $X$ is the circle bundle inside the dual of the quantizing line bundle, see Remark 1.1. A Hamiltonian and holomorphic action of a compact Lie group $G$ on $M$ produces an infinitesimal action of its Lie algebra on $X$. Under suitable topological conditions, for example if the group is semisimple or the manifold $M$ is simply connected, this action can be integrated to an action of $G$ on the circle bundle $X$ inside the dual of the line one, $L^{*}$. This action of $G$ on $X$ preserves $\omega_{0}$ and $J$.
Now, let $R$ be the Reeb vector field of the contact form $\omega_{0}$ and let $\mathcal{L}$ denote the Lie derivative. A vector field $V\,:\,X\rightarrow TX$ such that $\mathcal{L}_{V}\omega_{0}=g\,\omega_{0}$ for some smooth function $g\,:\,X\rightarrow\mathbb{R}$ is called a contact vector field.
By [MDS, Lemma 3.5.14], for every function $g\,:\,X\rightarrow\mathbb{R}$ there exists a unique contact vector field $X_{g}$ which satisfies
$\omega_{0}(X_{g})=g$
and
$\mathrm{d}\omega_{0}(X_{g},\,V)=(\mathrm{d}g(R))\,\alpha(V)-\mathrm{d}g(V)$
for each vector field $V$ on $X$. Thus, there is a one-to-one correspondence between contact vector fields and smooth functions on $(X,\,\omega_{0})$.
The analog of the Poisson bracket in contact geometry is
$\{g,\,h\}:=\omega_{0}([X_{g},\,X_{h}])=\mathrm{d}g(X_{h})-\mathrm{d}h(X_{g})+\mathrm{d}\omega_{0}(X_{g},\,X_{h})\,$
for $f$ and $g$ smooth function on $X$. Now, we are interested in the subspace of $H$-invariant smooth functions on $X$. Thus, for any $S^{1}$-invariant smooth function $g$ we have $\mathrm{d}g(R)=R(g)=0$ and
$\mathrm{d}\omega_{0}(X_{g},\,V)=-\mathrm{d}g(V)\,.$
Eventually, we note that for two $H$-invariant smooth functions $g$ and $h$ on $X$, we obtain
$\{g,\,h\}=-\mathrm{d}\omega_{0}(X_{g},\,X_{h})+\mathrm{d}\omega_{0}(X_{h},\,X_{g})+\mathrm{d}\omega_{0}(X_{g},\,X_{h})=\mathrm{d}\omega_{0}(X_{h},\,X_{g})\,.$
The CR reduction is the CR manifold $X_{G}:=\mu^{-1}(0)/G$, [HH]. Let $\pi_{X}\,:\,\mu^{-1}(0)\rightarrow X_{G}$ be the projection. Then, $X_{G}$ is a circle bundle over the symplectic reduction $M_{G}:=\Phi^{-1}(0)/G$. In Section 4 we explain how to generalize some of the results in [GS] to the setting of pseudo-Riemannian manifolds. In particular there is an identification
$M_{G}\cong M_{s}/G^{\mathbb{C}}$
where $G^{\mathbb{C}}$ is the complexification of the group $G$ and $M_{s}$ is the saturation by $G^{\mathbb{C}}$ of $\Phi^{-1}(0)$ in $M$. Furthermore, we shall extend the theorem “quantization commutes with reduction” to the pseudo-Riemannian setting. This theorem is well-known and has been proved in different contexts; we refer to [Ma] for a survey. We have decided to include this result here because we explain how to generalize some results along the way of the proof, which may be of interest for further study, see Section 4.
Now, consider a smooth function $\widetilde{f}$ on $X$ invariant under the action of $G$ and $S^{1}$, it gives rise to a smooth function $\widetilde{f}_{|Y}$ on $Y=\mu^{-1}(0)$. Then, $\widetilde{f}_{|Y}$ defines a smooth function $f$ on $M_{G}$. Thus, we define the sup-norm of a $G$-invariant function on $\Phi^{-1}(0)$ to be
$\lVert\widetilde{f}\rVert_{\infty}:=\lVert f\rVert_{\infty}:=\sup\left\{\lvert\widetilde{f}(x)\rvert\,:\,x\in\mu^{-1}(0)\right\}$ (5)
which is the uniform norm of $f$ on the quotient $M_{G}$. Vice-versa, let $f$ be a smooth $S^{1}$ invariant function on $X_{G}$, its lift defines a smooth $G$ and $S^{1}$ invariant function on $Y$. By [Lee, Lemma 5.34 (b)], the extension Lemma for functions on submanifolds, since $Y$ is a smooth properly embedded submanifold of $X$, then there exists a smooth function $\widetilde{f}$ on $X$ whose restriction to $Y$ coincides with $f$.
As the last piece of notation, let $\widetilde{f}$ be a $G$-invariant function on $X$, we denote the operator norm of its corresponding level $k$ $G$-invariant Toeplitz operator:
$\left\lVert T^{G}_{k}[\widetilde{f}]\right\rVert=\sup\left\{\frac{\lVert T^{G}_{k}[\widetilde{f}]s\rVert}{\lVert s\rVert}\,:\,s\in H^{G}_{k}(X)\,,s\neq 0\right\}\,.$ (6)
In the following theorem we state an analogue of the main result in [BMS]. Parts (c) and (b) of the following theorem are a consequence of the results contained in [GH1], see also [MM2] and [MM1] where the Bergman kernels and the corresponding Toeplitz operators on symplectic manifolds, for the mixed curvature case, are studied. Here, we prove Part (a) by adapting the strategy of [S2] to our setting.
Theorem 1.1.
Under the assumptions and notations above. Let $\widetilde{f},\,\widetilde{g}$ be $G$ and $S^{1}$ invariant functions on $X$, then
  • (a)
    there exists a positive constant $C$, such that
    $\lVert{f}\rVert_{\infty}-\frac{C}{k}\leq\left\lVert T^{G}_{k}[\widetilde{f}]\right\rVert\leq\lVert{f}\rVert_{\infty}\,$
    and in particular
    $\lim_{k}\left\lVert T^{G}_{k}[\widetilde{f}]\right\rVert=\lVert{f}\rVert_{\infty}\,;$
  • (b)
    the semi-classical Dirac condition holds:
    $\left\lVert ki\left[T^{G}_{k}[\widetilde{f}],\,T^{G}_{k}[\widetilde{g}]\right]-T^{G}_{k}[\{\widetilde{f},\,\widetilde{g}\}]\right\rVert=O\left(\frac{1}{k}\right)\,;$
  • (c)
    the semi-classical von Neumann’s condition holds:
    $\left\lVert T^{G}_{k}[\widetilde{f}]\,T^{G}_{k}\left[\widetilde{g}\right]-T^{G}_{k}\left[\widetilde{f}\cdot\widetilde{g}\right]\right\rVert=O\left(\frac{1}{k}\right)\,.$
Furthermore, the Berezin-Toeplitz quantization is also a positive quantization in the following sense, see [LeF]. Let $f\in C^{\infty}(X_{G})^{S^{1}}$ be such that $f(x)\geq 0$ for every $x\in X_{G}$. Then $T_{k}^{G}[f]$ is a non-negative operator, in the sense that
$\left(T_{k}^{G}[f]s,\,s\right)\geq 0$
for every $s\in H_{k}^{G}(X)$. Equivalently, by the spectral theorem, a self-adjoint linear operator $A$ is positive if and only if $\mathrm{Spec}(A)\subset[0,\,+\infty)$. On the other hand, the Weyl quantization of a non-negative function need not to be a positive function.
Definition 1.2.
A star product for the algebra $C^{\infty}(X_{G})^{S^{1}}$ is given by the power series
$g\star h=\sum_{j=0}^{+\infty}C_{j}(g,h)\,\nu^{-j}$
such that $\star$ is an associative $\mathbb{C}[[\nu]]$-linear product, that is,
$(g\star h)\star l=g\star(h\star l)\,,$
for all $g,h,l\in C^{\infty}(X_{G})^{S^{1}}$ and
$C_{0}(g,h)=g\cdot h\,,$
$C_{1}(g,h)-C_{1}(g,h)=i\,\{g,h\}\,,$
for all $g,h\in C^{\infty}(X_{G})^{S^{1}}$. The star product is said to be of Wick type if and only if one function is differentiated in holomorphic directions while the other is differentiated in anti-holomorphic directions only.
It is an easy consequence of [GH1] that the star product defined by Toeplitz operators is of Wick type. Our result prove that $\star$ defines a strict deformation quantization in the sense of [L, Chapter 2]. In the following Corollary we state the properties of the star product induced by Theorem 3.1. Here, we introduce a deformation quantization for $S^{1}$-invariant functions on $X_{G}$ induced by $G$ and $S^{1}$-invariant Toeplitz operators on $X$.
Corollary 1.1.
There exists a unique formal star-product
$f\star g:=\sum_{j=0}^{+\infty}\nu^{j}\,C_{j}(f,\,g)\,,\quad C_{j}(f,\,g)\in C^{\infty}(X_{G})^{S^{1}}$
such that for every smooth functions $f$ and $g$ on $X_{G}$ invariant under the circle action, for every $N\in\mathbb{N}$ we have with suitable constants $K_{N}(f,\,g)$, and for all $k\in\mathbb{N}$
$\left\lVert T^{G}_{k}[f]\,T^{G}_{k}[g]-\sum_{j=0}^{N-1}k^{-j}\,T^{G}_{k}[C_{j}(f,g)]\right\rVert=K_{N}(f,\,g)\,k^{-N}\,.$
Furthermore the star product $\star$ has the following properties.
  • (Unit)
    Let $1$ be the constant function equals to $1$ everywhere, then
    $1\star f=f\star 1=f$
    for every circle invariant smooth function on $X_{G}$.
  • (Parity)
    For every circle invariant smooth functions $f$ and $g$ on $X_{G}$, then
    $\overline{f\star g}=\overline{f}\star\overline{g}\,.$
  • (Trace)
    Denote the trace on $H(X)^{G}_{k}$ by $\mathrm{Tr}_{k}$. It induces a $\mathbb{C}[[\nu]]$-linear map
    $\mathrm{Tr}\,:\,{C}^{\infty}(X_{G})^{S^{1}}[[\nu]]\rightarrow\nu^{-n}\,\mathbb{C}[[\nu]]$
    such that
    $\mathrm{Tr}(f\star g)=\mathrm{Tr}(g\star f)\,.$
Now, if we put $G=\{e\}$, where $e$ is the identity element, we recover previous results in [S1], and we generalize it to pesudo-Kähler manifolds.
2 Preliminaries
2.1 Standard notations
The following notations are adopted through this article: $\mathbb{N}$ is the set of natural numbers, $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$, $\mathbb{R}$ is the set of real numbers, and we denote by +=${\mathbb{R}}_{+}=\{x\in\mathbb{R};\,x>0\}$ and ¯+=$\overline{\mathbb{R}}_{+}=\{x\in\mathbb{R};\,x\geq 0\}$. Furthermore, we write $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{N}^{n}_{0}$ if $\alpha_{j}\in\mathbb{N}_{0}$, $j=1,\ldots,n$.
[…]
Let $Y\subset M$ be an open set, the spaces of distribution sections of $B$ over $Y$ and smooth sections of $B$ over $Y$ will be denoted by $D^{\prime}(Y,B)$ and ${C}^{\infty}(Y,B)$, respectively. Let $E^{\prime}(Y,B)$ be the subspace of $D^{\prime}(Y,B)$ whose elements have compact support in $Y$.
Now, we recall the Schwartz kernel theorem. Let $B$ and $E$ be smooth vector bundles over paracompact orientable smooth manifolds $M$ and $N$, respectively, equipped with smooth densities of integration. If
$A:{C}^{\infty}_{0}(N,\,E)\rightarrow D^{\prime}(M,\,B)$
is continuous, we write $A(x,y)$ to denote the distribution kernel of $A$. The following two statements are equivalent
  1. 1.
    $A$ is continuous: $E^{\prime}(N,\,E)\rightarrow{C}^{\infty}(M,\,B)$,
  2. 2.
    The distributional kernel $A(x,y)\in{C}^{\infty}(M\times N,B\boxtimes E^{*})$.
If $A$ satisfies (1) or (2), we say that $A$ is smoothing on $M\times N$. Let
$A,\hat{A}:{C}^{\infty}_{0}(N,\,E)\rightarrow{D}^{\prime}(M,\,B)$
be continuous operators, then we write
$A\equiv\hat{A}$ (on $M\times N$)
if $A-\hat{A}$ is a smoothing operator. If $M=N$, we simply write “on $M$”. We say that $A$ is properly supported if the restrictions of the two projections $(x,y)\mapsto x$, $(x,y)\mapsto y$ to ${\rm supp\,}(A(x,y))$ are proper.
Eventually, let
$H(x,y)\in D^{\prime}(M\times N,\,B\boxtimes E^{*})\,$
then we write $H$ to denote the unique continuous operator
$C^{\infty}_{0}(N,E)\rightarrow D^{\prime}(M,B)$
with distribution kernel $H(x,y)$ and we identify $H$ with $H(x,y)$.
2.2 Operators and symbols
Let $(X,T^{1,0}X)$ be a compact, connected and orientable CR manifold of dimension $12n+$ 1, $n\geq 1$, we adopt the same notation as in Introduction.
Let $D$ be an open set of $X$. Recall that $\Omega^{0,q}(D)$ denote the space of smooth sections of $T^{*0,q}X$ over $D$ and let $\Omega^{0,q}_{c}(D)$ be the subspace of $\Omega^{0,q}(D)$ whose elements have compact support in $D$. Let
¯$\overline{\partial}_{b}:\Omega^{0,q}(X)\rightarrow\Omega^{0,q+1}(X)$
[…]
¯$\overline{\partial}_{b}:{\rm Dom\,}\overline{\partial}_{b}\subset L^{2}_{(0,r)}(X)\rightarrow L^{2}_{(0,r+1)}(X)\,,$
where
${\rm Dom\,}\overline{\partial}_{b}:=\{u\in L^{2}_{(0,r)}(X);\,\overline{\partial}_{b}u\in L^{2}_{(0,r+1)}(X)\}$
and, for any $u\in L^{2}_{(0,r)}(X)$, ¯$\overline{\partial}_{b}u$ is defined in the sense of distributions. We also write
¯$\overline{\partial}^{*}_{b}:{\rm Dom\,}\overline{\partial}^{*}_{b}\subset L^{2}_{(0,r+1)}(X)\rightarrow L^{2}_{(0,r)}(X)$
[…]
$\Big{\{}s\in L^{2}_{(0,q)}(X);\,s\in{\rm Dom\,}\overline{\partial}_{b}\cap{\rm Dom\,}\overline{\partial}^{*}_{b},\,\overline{\partial}_{b}s\in{\rm Dom\,}\overline{\partial}^{*}_{b},\,\overline{\partial}^{*}_{b}s\in{\rm Dom\,}\overline{\partial}_{b}\Big{\}}$
and
$\Box^{q}_{b}s=\overline{\partial}_{b}\overline{\partial}^{*}_{b}s+\overline{\partial}^{*}_{b}\overline{\partial}_{b}s\>\>\text{for $s\in{\rm Dom\,}\Box^{q}_{b}$}\,.$
Let
$S^{(q)}:L^{2}_{(0,q)}(X)\rightarrow{\rm Ker\,}\Box^{q}_{b}$ (7)
[…]
$S^{(q)}(x,y)\in{D}^{\prime}(X\times X,T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
denote the distribution kernel of $S^{(q)}$.
Now, we recall Hörmander symbol space. Let $D\subset X$ be a local coordinate patch with local coordinates $x=(x_{1},\ldots,x_{2n+1})$.
Definition 2.1.
For $m\in\mathbb{R}$,
$S^{m}_{1,0}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
is the space of all
$a(x,y,t)\in{C}^{\infty}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
such that, for all compact $K\Subset D\times D$ and all $\alpha,\beta\in\mathbb{N}^{2n+1}_{0}$, $\gamma\in\mathbb{N}_{0}$, there is a constant $C_{\alpha,\beta,\gamma}>0$ such that
$\left|\partial^{\alpha}_{x}\partial^{\beta}_{y}\partial^{\gamma}_{t}a(x,y,t)\right|\leq C_{\alpha,\beta,\gamma}(1+\left|t\right|)^{m-\gamma},\ \ \mbox{forevery $(x,y,t)\in K\times\mathbb{R}_{+},t\geq 1$}.$
Put
$S^{-\infty}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
to be
$\bigcap_{m\in\mathbb{R}}S^{m}_{1,0}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*}).$
Now, let
$a_{j}\in S^{m_{j}}_{1,0}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})\,,$
$j=0,1,2,\ldots$ with $m_{j}\rightarrow-\infty$, as $j\rightarrow\infty$. Then there exists
$a\in S^{m_{0}}_{1,0}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
unique modulo $S^{-\infty}$, such that
$a-\sum^{k-1}_{j=0}a_{j}\in S^{m_{k}}_{1,0}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*}\big{)}$
for $k=0,1,2,\ldots$.
If $a$ and $a_{j}$ have the properties above, we write $a\sim\sum^{\infty}_{j=0}a_{j}$ in
$S^{m_{0}}_{1,0}\big{(}D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*}\big{)}\,.$
Furthermore, we write
$s(x,y,t)\in S^{m}_{{\rm cl\,}}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
if
$s(x,y,t)\in S^{m}_{1,0}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
and
$s(x,y,t)\sim\sum^{\infty}_{j=0}s^{j}(x,y)t^{m-j}\text{ in }S^{m}_{1,0}(D\times D\times\mathbb{R}_{+}\,,T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})\,,$
where
$s^{j}(x,y)\in{C}^{\infty}(D\times D,T^{*0,q}X\boxtimes(T^{*0,q}X)^{*}),\ j\in\mathbb{N}_{0}\,.$
Sometimes we simply write $S^{m}_{1,0}$ to denote
$S^{m}_{1,0}(D\times D\times\mathbb{R}_{+},T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})\,,$
where $m\in\mathbb{R}\cup\left\{-\infty\right\}$.
Let $E$ be a smooth vector bundle over $X$. Let $m\in\mathbb{R}$, $0\leq\rho,\delta\leq 1$. Let $S^{m}_{\rho,\delta}(X,E)$ denote the Hörmander symbol space on $X$ with values in $E$ of order $m$ type $(\rho,\delta)$ and let $S^{m}_{{\rm cl\,}}(X,E)$ denote the space of classical symbols on $X$ with values in $E$ of order $m$. For $a\in S^{m}_{{\rm cl\,}}(X,E)$, we write ${\rm Op\,}(a)$ to denote the pseudodifferential operator on $X$ of order $m$ from sections of $E$ to sections of $E$ with full symbol $a$.
Let $D\subset X$ be an open set. Let
$L^{m}_{{\rm cl\,}}(D,T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
denote the space of classical pseudodifferential operators on $D$ of order $m$ from sections of $T^{*0,q}X$ to sections of $T^{*0,q}X$ respectively. Let
$P\in L^{m}_{{\rm cl\,}}(D,T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})\,$
and we write $\sigma_{P}$ and $\sigma^{0}_{P}$ to denote the full symbol of $P$ and the principal symbol of $P$ respectively.
2.3 Toeplitz operators
In this sections we briefly recall some results from [GH1]. First let us recall the assumptions from Introduction. Here, the Reeb vector field $R$ is transversal to the space $\underline{\mathfrak{g}}$ at every point $p\in\mu^{-1}(0)$,
$e^{i\theta}\cdot g\circ x=g\circ e^{i\theta}\cdot x,\ \ \mbox{for every $x\in X$, $\theta\in[0,2\pi[$, $g\in G$,}$
and
The action of $G$ and $S^{1}$ is free near $\mu^{-1}(0)$.
For every smooth function $g\in{C}^{\infty}(X)^{G}$ invariant under the action of $S^{1}$, the $k$-th $G$-invariant Toeplitz operator is given by
$T^{G}_{k}[g]:=S^{G}_{k}\circ g\cdot\circ S^{G}_{k}:L^{2}_{(0,n_{-})}(X)\rightarrow L^{2}_{(0,n_{-}),k}(X)^{G}.$ (8)
The main interest for us is the case when $g=\tilde{f}$ and $f$ is in $C^{\infty}(X_{G})^{S^{1}}$. For the following theorem see Theorem $1.8$ in [HH] and [GH1].
Theorem 2.1.
Let $f\in{C}^{\infty}(X)^{G}$ and suppose $q=n_{-}$. Let $D$ be an open set in $X$ such that the intersection $\mu^{-1}(0)\cap D=\emptyset$. Then,
$T^{G}_{k}[f]\equiv O(k^{-\infty})$
on $D$.
Let $p\in\mu^{-1}(0)$ and let $U$ a local neighborhood of $p$. Then, if $q=n_{-}$, for every fixed $y\in U$, we consider $T_{k}^{G}[f](x,y)$ as a $k$-dependent smooth function in $x$. There exist local coordinates $(x_{1},\dots\,x_{2n+1})$ on $U$ such that
$T^{G}_{k}[f](x,\,y)=e^{ik\,\Psi(x,\,y)}\,b(x,\,y,\,k)+O(k^{-\infty})\,$
for every $x,\,y\in U$. For an explicit expression of the phase function $\Psi$ we refer to [GH1], the symbol satisfies
$b(x,\,y,\,m)\in S^{n-d/2}_{\mathrm{loc}}(1,\,U\times U,\,T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})\,$
and the leading term of $b(x,y,m)$ along the diagonal at $x\in\mu^{-1}(0)$ is given by
$b_{0}(x,\,x):=2^{d-1}\frac{1}{V_{{\rm eff\,}}(x)}\pi^{-n-1+\frac{d}{2}}\left|\det R_{x}\right|^{-\frac{1}{2}}\left|\det{L}_{x}\right|\tau_{x,n_{-}}$
where
$V_{{\rm eff\,}}(x):=\int_{O_{x}}\mathrm{dV}_{O_{x}}\,,$
$O_{x}$ is the orbit through $x$, and $R_{x}$ is the linear map
$\begin{array}[]{rll}R_{x}:\underline{\mathfrak{g}}_{x}&\rightarrow&\underline{\mathfrak{g}}_{x},\\ u&\rightarrow&R_{x}u,\ \ \langle\,R_{x}u\,|\,v\,\rangle=\langle\,\mathrm{d}\omega_{0}(x)\,,\,Ju\wedge v\,\rangle.\end{array}$
Now we recall the definition of the operator $\hat{H}$, which appears in the description of the $C_{j}$’s, see [GH1]. It is convent to introduce a class of operator which looks microlocally like the Szegő kernel. Here we recall the definition from [GH1] for $q=n_{-}$ and action free case.
Definition 2.2.
Let $H:\Omega^{0,n_{-}}(X)\rightarrow\Omega^{0,n_{-}}(X)$ be a continuous operator with distribution kernel
$H(x,y)\in{D}^{\prime}(X\times X,T^{*0,n_{-}}X\boxtimes(T^{*0,n_{-}}X)^{*})\,.$
We say that $H$ is a complex Fourier integral operator of $G$-invariant Szegő type of order $k\in\mathbb{Z}$ if $H$ is smoothing away $\mu^{-1}(0)$ and $H$ has the following microlocal expression. Let $p\in\mu^{-1}(0)$, and let $D$ a local neighborhood of $p$ with local coordinates $(x_{1},\dots\,x_{2n+1})$. Then, the distributional kernel of $H$ satisfies
$H(x,y)\equiv\int^{\infty}_{0}e^{it\,\Phi_{-}(x,y)}a_{-}(x,\,y,\,t)\,\mathrm{d}t$
on $D$ where
$a_{-}\in S^{k-\frac{d}{2}}_{{\rm cl\,}}(D\times D\times\mathbb{R}_{+},T^{*0,n_{-}}X\boxtimes(T^{*0,q}X)^{*})\,,$
$a_{+}=0$ if $q\neq n_{+}$, where $\Phi_{-}$ is the phase of the $G$-invariant Szegő kernel. Furthermore, we write $\sigma^{0}_{H,-}(x,y)$ to denote the leading term of the expansion of $a_{-}$.
Let $\Psi_{k}(X)^{G}$ denote the space of all complex Fourier integral operators of Szegő type of order $k$.
Here we adapt the discussion before Theorem 4.9 in [GH1] to $G$-action case. Let
$\hat{R}:=\frac{1}{2}S^{G}(-iR+(-iR)^{*})S^{G}:\Omega^{0,n_{-}}(X)\rightarrow\Omega^{0,n_{-}}(X),$ (9)
[…]
$\begin{split}&S^{G}\hat{H}=\hat{H}=\hat{H}S^{G}\,,\\ &\hat{R}\hat{H}\equiv\hat{H}\hat{R}\equiv S^{G}\,.\end{split}$ (10)
Note that $\hat{H}$ is uniquely determined by (10), up to some smoothing operators. The operator $\hat{H}$ can always be found by setting
$\hat{H}=S^{G}\tilde{H}S^{G}$
for a given pseudodifferential operator $\tilde{H}$. In fact, the first equation in (10) is always satisfied. Furthermore the second equation can be solved by an easy application of the stationary phase formula.
Now, we can recall the $G$-invariant version of Theorem 4.9 in [GH1].
Theorem 2.2.
Recall that we work with the assumption that $q=n_{-}$. Let $f,g\in{C}^{\infty}(X)$, then we have
$T^{G}[f]\circ T^{G}[g]-\sum^{N}_{j=0}\hat{H}^{j}\,T^{G}\left[{\hat{C}_{j}(f,g)}\right]\in\Psi_{n-N-1}^{G}(X),$
for every $N\in\mathbb{N}_{0}$, where $\hat{C}_{j}(f,g)\in{C}^{\infty}(X)$, $\hat{C}_{j}$ is a universal bidifferential operator of order $\leq 2j$, $j=0,1,\ldots$, and
$\begin{split}\hat{C}_{0}(f,g)&=f\cdot g,\\ \hat{C}_{1}(f,g)-\hat{C}_{1}(g,f)&=i\,\{f,g\}.\end{split}$
Eventually, we recall the following theorem from [GH1].
Theorem 2.3.
With the same assumptions as above, let $f,g\in{C}^{\infty}(X)^{G}$. Let $q=n_{-}$. Then, as $k\gg 1$,
$\left\|T^{G}_{f,k}\circ T^{G}_{g,k}-T^{G}_{g,k}\circ T^{G}_{f,k}-\frac{1}{k}\,T^{G}_{i\{f,g\},k}\right\|=O(k^{-2}),$ (11)
and
$\left\|T^{G}_{f,k}\circ T^{G}_{g,k}-\sum^{N}_{j=0}k^{-j}\,T^{G}_{C_{j}(f,g),k}\right\|=O(k^{-N-1})$ (12)
in $L^{2}$ operator norm, for every $N\in\mathbb{N}$, where $C_{j}(f,g)\in{C}^{\infty}(X)^{G}$, $C_{j}$ is a universal bidifferential operator of order $\leq 2j$, $j=0,1,\ldots$, and
$\begin{split}C_{0}(f,g)&=f\cdot g\,,\\ C_{1}(f,g)-C_{1}(g,f)&=i\,\{f,g\}\,.\end{split}$ (13)
Moreover, the star product
$f*g=\sum_{j=0}^{+\infty}C_{j}(f,g)\nu^{-j},$ (14)
$f,g\in{C}^{\infty}(X)^{G}_{0}$, is associative.
3 Berezin transform
3.1 Vector valued reproducing kernel
Here, we recall some definitions concerning reproducing kernel Hilbert spaces for vector valued functions, we refer to [BM] and [Pe] (see also [CVT, Section 2]).
Definition 3.1.
Let $X$ be a manifold and $E$ be a vector bundle on $X$. A $E$-valued reproducing kernel Hilbert space on $X$ is a Hilbert space $H$ such that:
  • (1)
    the elements of $H$ are sections of the bundle $E\rightarrow X$,
  • (2)
    for all $x\in X$ there exists a positive constant $C_{x}$ such that
    $\lVert s(x)\rVert_{E_{x}}\leq C_{x}\lVert s\rVert_{H}\,,$
    for each $s\in H$.
[…]
$\mathrm{ev}_{x}\,:\,H^{G}_{k}(X)\rightarrow T^{*\,(0,q)}X,\qquad\mathrm{ev}_{x}(s)=s(x)$
is a bounded operator and the reproducing kernel associated with $H^{G}_{k}(X)$ is $S^{G}$. The kernel $S^{G}$ reproduces the value of a section $s\in H^{G}_{k}(X)$ at point $x\in X$. Indeed, for all $x\in X$ and $v\in T^{*\,(0,q)}_{x}X$
$\mathrm{ev}_{x}^{*}\,v=S^{G}_{k}(\cdot,\,x)\,v$
so that
$\langle s(x)|v\rangle=(s,\,S^{G}_{k}(\cdot,\,x)v)\,.$
Thus, for each $x\in X$ and $v\in T^{*\,(0,q)}_{x}X$ the $(0,q)$-forms $s_{x,v}=S^{G}_{k}(\cdot,\,x)\,v$ is called the coherent state (of level $k$) associated to $x$ and $v$.
Definition 3.2.
The covariant Berezin symbol $\sigma^{(k)}_{k}(A)$ of an equivariant operator of Szegő type $A\in\Psi^{G}_{k}(X)$ is
$\sigma_{v}^{(k)}(A)\,:\,X\rightarrow\mathbb{C},\qquad x\mapsto\sigma_{v}(A)(x):=\frac{\left(s_{x,v}|\,As_{x,v}\right)}{\left(s_{x,v}|\,s_{x,v}\right)}$
for a given fixed $v\in T^{*(0,q)}_{x}X$.
Remark 3.1.
The definition of coherent states goes back to Berezin. A coordinate independent version and extensions to line bundles were given by Rawnsley [R], we mainly refer to [S2] where the coherent vectors are parameterized by the elements of $L^{*}\setminus 0$, where $L$ is a positive line bundle of a given Hodge manifold. In the following we shall generalize the coherent states in our setting.
Let $f$ be a $G$ invariant function on $\mu^{-1}(0)$, for any extension $\widetilde{f}$ on $X$, we can assign to it its Toeplitz operator $T^{G}_{k}[\widetilde{f}]$ and then assign to it the covariant symbol, as in Definition 3.2
$\sigma_{v}^{(k)}\left(T^{G}_{k}[\widetilde{f}]\right)\,:\,X\rightarrow\mathbb{C}$
which is again an element of $C^{\infty}(X)^{G}$ and thus it defines a circle invariant function on the CR reduction $X_{G}$.
3.2 Berezin transforms for Toeplitz operators
Recall that $\pi\,:\,\mu^{-1}(0)\rightarrow X_{G}$ is the projection. Given a smooth function $f\in C^{\infty}(X_{G})$ invariant under the circle action, we define:
Definition 3.3.
The map
$C^{\infty}(X_{G})^{S^{1}}\rightarrow C^{\infty}(X_{G})^{S^{1}}\,,\qquad f\mapsto I_{k}^{G}[f]:=\sigma^{(k)}_{v}(T^{G}_{k}[f])$
is called the Berezin transform of level $k$.
In particular, we are interested in studying $I_{k}^{G}[f]$ on a neighborhood $U$ of $p$ in $\mu^{-1}(0)$.
Theorem 3.1.
Let $p\in\mu^{-1}(0)$ and let $U$ a local neighborhood of $p$. Then, if $q=n_{-}$, there exist local coordinates $(x_{1},\dots\,x_{2n+1})$ on $U$ such that the Berezin transform $I^{G}_{k}[f]$ evaluated at the point $x\in U$ has a complete asymptotic expansion in decreasing integer powers of $k$
$I^{G}_{k}[f](x)\sim\sum_{j=0}^{+\infty}I_{j}[f](x)\,\frac{1}{k^{j}}\,,$
as $k$ goes to infinity, where $I_{j}\,:\,C^{\infty}(X_{G})^{S^{1}}\rightarrow C^{\infty}(X_{G})^{S^{1}}$ are maps such that $I_{0}[f]=f$.
Proof of Theorem 3.1.
Let $f\in C^{\infty}(X_{G})^{S^{1}}$ and let us first study the numerator of
$\sigma^{(k)}_{v}(T^{G}_{k}[\widetilde{f}])$
which is
$\left(S^{G}_{k}(\cdot,\,x)\,v\bigg{|}T^{G}_{k}[\widetilde{f}]\,S^{(q)}_{k}(\cdot,\,x)\,v\right)=\int_{X}v^{\dagger}\,S^{(q)}_{k}(x,\,y)\widetilde{f}(y)\,S^{(q)}_{k}(y,\,x)\,v\,\mathrm{dV}_{X}(y)\,.$
The denominator is obtain by taking $f\equiv 1$. Thus, by Proposition 2.1, we have a complete asymptotic expansion in decreasing integer power of $k$, and we obtain
$\left(S^{(q)}_{k}(\cdot,\,x)\,v\bigg{|}T^{(q)}_{k}[\widetilde{f}]\,S^{(q)}_{k}(\cdot,\,x)\,v\right)\sim k^{d}\,v^{\dagger}\sigma^{0}_{T^{G}[\widetilde{f}]}(x,\,x)\,v+O(k^{d-1})$
where
$\sigma^{0}_{T^{G}[\widetilde{f}]}(x,x)=\frac{1}{2}\pi^{-n-1}\left|{\rm det\,}L_{x}\right|\,\widetilde{f}(x)\,\tau_{x,n_{-}}\,,$
for every $x\in\mu^{-1}(0)$. Thus, taking the quotient we get
$I_{k}^{G}[\widetilde{f}](x)=\widetilde{f}(x)+O(k^{d-1})\,$
for each $x\in\mu^{-1}(0)$. ∎
Theorem 3.1 is consequence of the asymptotic expansion of Toeplitz operators for $(0,q)$-forms and the stationary phase formula. Here, we shall mainly refer to [GH1] where we study Toeplitz operators on CR manifolds, see also [MM2] and [Hs] for the asymptotic behavior of the Bergman kernel for $(0,q)$-forms, and references therein.
4 Quotients and complexification of groups
Here, we recall some results proved in [GS, Section 4] and we explain how to extend it to pseudo Kähler manifolds. There exists a unique connected complex Lie group, $G^{\mathbb{C}}$, with the following two properties:
  • its Lie algebra is $\mathfrak{g}\oplus\sqrt{-1}\,\mathfrak{g}$,
  • $G$ is a maximal compact subgroup of $G^{\mathbb{C}}$.
Now, we shall prove some properties of $G$-actions on pseudo-Kähler manifolds. Now let $M$ be a $2n$-dimensional symplectic manifold. A polarization of $X$ is an integrable Lagrangian subbundle, $F$, of $TM\otimes\mathbb{C}$. A symplectic manifold is a pseudo-Kähler manifold if it possesses a polarization $F$. The next lemma, an analogue of [GS, Lemma 4.3], will be used to reconcile this definition with the standard one:
Lemma 4.1.
Let $V$ be a (real) symplectic vector space with symplectic form, $\Omega$. Let $F$ be a Lagrangian subspace of $V\otimes\mathbb{C}$ Then there exists a unique linear mapping $J\,:\,V\rightarrow V$ such that
  • i)
    $J^{2}=-I$,
  • ii)
    $F=\{v+\sqrt{-1}\,Jv,\,v\in V\}$,
  • iii)
    $\Omega(Jv,\,Jw)=\Omega(v,\,w)$,
  • iv)
    The quadratic form
    $B(v,\,w)=\Omega(v,Jw)$
    is symmetric and non-degenerate.
Let $M$ be a symplectic manifold and $F$ a polarization. By the lemma we get for each $m\in M$ a mapping $J_{m}\,:\,T_{m}M\rightarrow T_{m}M$ with the properties i), ii), and iii) and a quadratic form, $B_{m}$, on $M$. Furthermore $J$ defines an almost-complex structure on $M$ and $B$ a pseudo-Riemannian structure. The integrability of $F$ implies that the almost-complex structure is complex. Therefore, the quadruple $(M,J,B,\Omega)$ is pseudo-Kähler manifold in the usual sense.
Let $(M,F)$ be a compact Kähler manifold and $G$ a compact connected Lie group which acts on $M$, preserving $F$. We shall prove Theorem 4.1 below, which is an analogue of [GS, Theorem 4.4] and in fact the first part of the proof is identical. However, note that [GS, Theorem 4.4] relies on the fact that the group of analytic diffeomorphisms of $M$ which preserve $F$ is a finite dimensional Lie group, see [K]. Here, we prove that the former result still holds true in the setting of pseudo-Riemannian manifolds (in Theorem 4.1 we assume that $G^{\mathbb{C}}$ is simply connected, for the general case we refer to [GS, Theorem 4.4]).
Let $\mathrm{Iso}(X,\,g)$ be the group of isometries from $(M,\,g)$ onto itself, that is $f\in\mathrm{Iso}(X,\,g)$ if and only if $f$ is a smooth diffeomorphism and $f^{*}g=g$. Let $(M,\,T^{1,0}M)$ be a connected complex manifold and denote by $\mathrm{Iso}(M,\,g)$ the group of isometries on $M$ with respect to some Riemannian metric $g$. Let $\mathrm{Aut}_{J}(M)$ be the group of complex automorphisms on $M$, that is $f\in\mathrm{Aut}_{J}(M)$ if and only if $f\,:\,M\rightarrow M$ is a smooth diffeomorphism satisfying $\mathrm{d}f(T^{1,0}M)\subseteq T^{1,0}M$.
Theorem 4.1.
Let us assume that $G^{\mathbb{C}}$ is simply-connected. The action of $G$ can be canonically extended to an action of $G^{\mathbb{C}}$, preserving the pseudo- Kähler structure of $M$.
Proof.
Let
$\tau\,:\,\mathfrak{g}^{\mathbb{C}}\rightarrow\text{(real) vector fields on $M$}$
be the mapping,
$\xi^{(1)}+\sqrt{-1}\,\xi^{(2)}\mapsto\xi^{(1)}_{M}+J\,\xi^{(2)}_{M}$
where $\xi^{(1)},\,\xi^{(2)}\in\mathfrak{g}$. By (4.2) and (4.3) in [GS], $\tau$ is a morphism of Lie algebras. Moreover, by (4.2), if $\eta\in\mathfrak{g}^{\mathbb{C}}$, $\tau(\eta)$ is a vector field preserving $F$. By Lemma 4.2 below, $\mathrm{Iso}(M,\,g)\cap\mathrm{Aut}_{J}(M)$ is a (finite dimensional) Lie group; therefore, if $G^{\mathbb{C}}$ is simply-connected, $\tau$ can be extended uniquely to a morphism of Lie groups. ∎
Lemma 4.2.
The group $\mathrm{Iso}(M,\,g)\cap\mathrm{Aut}_{J}(M)$ is a Lie group.
Proof.
By [Ph] and [K], we have that $\mathrm{Iso}(M,\,g)$ is a finite-dimensional Lie group, since $(M,\,g)$ is a pseudo-Riemannian manifold. Recall that a subgroup of a Lie group is itself a Lie group if it is a closed subgroup.
Now, consider the subgroup $\mathrm{Iso}(M,\,g,\,J)\subset\mathrm{Iso}(M,\,g)$, consisting of those diffeomorphisms that preserve both the metric $g$ and the complex structure $J$. The preservation of $g$ and $J$ by a diffeomorphism are smooth conditions. Thus, we can consider the set $\mathrm{Iso}(M,\,g,\,J)$ as the solution set of these equations within the group $\mathrm{Iso}(M,\,g)$. If a sequence of diffeomorphisms in $\mathrm{Iso}(M,\,g,\,J)$ converges to a diffeomorphism in the topology of $\mathrm{Iso}(M,\,g)$ (which is the compact-open topology), the limit will also preserve $g$ and $J$.
Thus, $\mathrm{Iso}(M,\,g,\,J)$ is a closed subgroup of the Lie group $\mathrm{Iso}(M,\,g)$, it is itself a Lie group. ∎
Recall now that $(M,\,\omega)$ is the complex pseudo-Kähler manifold obtained by quotienting out by the circle action on $X$, $M:=X/S^{1}$. The CR action of $G$ on $X$ descends to a complex and Hamiltonian action on $M$ with moment map $\Phi\,:\,M\rightarrow\mathfrak{g}^{*}$. Thus, we can form the reduced space
$M_{G}:=\Phi^{-1}(0)/G\,.$
Let $M_{s}$ be the saturation of $\Phi^{-1}(0)$ with respect to $G^{\mathbb{C}}$,
$M_{s}:=\{g\circ x\,:\,x\in\Phi^{-1}(0),\,g\in G^{\mathbb{C}}\}\,,$
the points of $M_{s}$ are called stable points. In the same way as in [GS, Theorem 4.5], one can prove the following theorem. Here, we shall retrace the proof again since the assumption (4) will be crucial in the pseudo-Riemannian case.
Theorem 4.2.
The set $M_{s}$ is an open subset of $M$ and $G^{\mathbb{C}}$ acts freely on it. Then, $M_{G}$ can be represented as the quotient space $M_{G}:=M_{s}/G^{\mathbb{C}}$.
Proof.
Let $V$ be a real symplectic vector space and $F$ a Lagrangian subspace of the complexification $V\otimes\mathbb{C}$. Let $J$ and $B$ as in Lemma 4.1.
Since the origin in $\mathfrak{g}$ is a regular value of $\Phi$, then $\Phi^{-1}(0)$ is a $G$-invariant co-isotropic submanifold of $M$. Moreover, the action of $G$ on $\Phi^{-1}(0)$ is locally free and the orbits of $G$ are the leaves of the null-foliation.
Recall that $b$ is the nondegenerate bilinear form on $HX$ such that
$b(\cdot,\cdot)=\mathrm{d}\omega_{0}(\cdot,J\cdot)\,.$
Let us denote $\mu^{-1}(0)$ by $Y$. There is a natural projection $\pi_{Y}\,:\,Y\rightarrow\Phi^{-1}(0)$. Recall that we assume that 00 is a regular value of $\mu$, and that the action of $G$ on $\mu^{-1}(0)$ is free and
[…]
where
$\underline{\mathfrak{g}}={\rm Span\,}(\xi_{X};\,\xi\in\mathfrak{g})$
and
$\underline{\mathfrak{g}}^{\perp_{b}}=\left\{v\in HX;\,b(\xi_{X},v)=0,\ \ \forall\xi_{X}\in\underline{\mathfrak{g}}\right\}\,.$
By [HH, Section 2.5], we observe that when restricted to $\mathfrak{g}\times\mathfrak{g}$, the bilinear form $b$ is nondegenerate on $Y$. Furthermore,
$\pi_{Y}\left(\underline{\mathfrak{g}}_{x}\right)=T_{\pi(x)}\Phi^{-1}(0)\,$
and
$\left\{{\xi}_{M}(\pi(x))\in T_{\pi(x)}M\,:\,\xi\in\sqrt{-1}\,\mathfrak{g}\right\}$
is a complementary space to $T_{\pi(x)}\Phi^{-1}(0)$ in $T_{\pi(x)}M$. This shows that $M_{s}$ contains an open neighborhood, $U$, of $\Phi^{-1}(0)$. Since
$M_{s}=\bigcup_{g\in G^{\mathbb{C}}}g\,U$
then $M_{s}$ is itself open. Thus, the stabilizer algebra of $x$ in $\mathfrak{g}^{\mathbb{C}}$ is zero; so the action of $G^{\mathbb{C}}$ on $M_{s}$ is locally free. To show that $G^{\mathbb{C}}$ acts freely on $M_{s}$ one can proceed as in the second part of [GS, Theorem 4.5] at page 527. ∎
Now, we shall need the following theorem.
Theorem 4.3.
The set $M\setminus M_{s}$ is contained in a complex sub-manifold of $M$ of complex codimension greater than
[…]
This theorem is an analogue of [GS, Theorem 5.7] for pseudo-Riemannian manifolds. Note that here the existence theorem [GS, Theorem 5.6] is not true in general for the pseudo-Riemannian case: we have to replace the spaces of smooth sections of $L^{k}$ with space of $L^{k}$-valued $(0,\,n_{-})$-forms, see Theorem 4.5. Then, Theorem 4.3 follows by combining Theorem 4.4 and 4.5 below.
Theorem 4.4.
Let $s$ be a $G$-invariant $L^{k}$-valued $(0,\,n_{-})$-forms on $M$. Then, given $m\in M$ if $s(m)\neq 0$, then $m\in M_{s}$.
Proof.
See the proofs of Theorems 5.3 and 5.4 in [GS], they hold true in our setting up to replacing the space of $G$-invariant holomorphic sections of $L^{k}$ with the space of $G$-invariant $L^{k}$-valued $(0,\,n_{-})$-forms. ∎
Theorem 4.5.
If the set $\Phi^{-1}(0)$ is non-empty and zero is a regular value of $\Phi$, then for some $k$, there exists a global non-vanishing holomorphic $G$-invariant $L^{k}$-valued $(0,\,n_{-})$-forms.
Proof.
This is a consequence of the asymptotic expansions of $G$-invariant Szegő kernel, see [HH]. Indeed note that
$\dim_{\mathbb{C}}H^{G}_{k}(X)=\int_{X}S^{G}_{k}(x,\,x)\,\mathrm{dV}_{X}(x)\,.$
Now, since $S^{G}_{k}$ has a non-trivial asymptotic expansion on $\Phi^{-1}(0)$ in decreasing integer power of $k$, whose leading term has degree $n$. Thus, for $k$ sufficiently large $H^{G}_{k}(X)$ can not be empty. ∎
Now, “quantization commutes with reduction” follows as in [GS] by combining [GS, Theorem 5.7] (the analogous result here is Theorem 4.3) and [GS, Theorem 5.2].
5 Proof of Theorem 1.1
Let $\widetilde{f}\in C^{\infty}(X)$ be a $G$ and $S^{1}$ smooth invariant function on $X$. First, we prove that
$\left\lVert I_{k}^{G}[f]\right\rVert_{\infty}\leq\rVert T^{G}_{k}[\widetilde{f}]\lVert\leq\lVert f\rVert_{\infty}\,.$ (15)
Recall that given two $G$ invariant $(0,\,n_{-})$ forms $s$ and $t$ in $H(X)^{G}_{k}$, we denote their inner product by $\left(s|t\right)$. Now, using the Cauchy-Schwartz inequality, we obtain that for each element $x\in\mu^{-1}(0)$
$\left\lVert I^{G}_{k}[f](x)\right\rVert^{2}_{\infty}$ $=\frac{\left|\left(S^{G}_{k}(\cdot,\,x)\,v\bigg{|}T^{G}_{k}[f]\,S^{G}_{k}(\cdot,\,x)\,v\right)\right|^{2}}{\left(S^{G}_{k}(\cdot,\,x)\,v\bigg{|}\,S^{G}_{k}(\cdot,\,x)\,v\right)^{2}}$
$\leq\frac{\left(T^{G}_{k}[f]\,S^{G}_{k}(\cdot,\,x)\,v\bigg{|}T^{G}_{k}[f]\,S^{G}_{k}(\cdot,\,x)\,v\right)}{\left(S^{G}_{k}(\cdot,\,x)\,v\bigg{|}\,S^{G}_{k}(\cdot,\,x)\,v\right)}$
$\leq\rVert T^{G}_{k}[f]\lVert^{2}$
where the last inequality follows by the definition of the operator norm, see (6). Taking the $\sup$ both sides over $\mu^{-1}(0)$, it shows the first inequality in (15).
For the second inequality, we note that for $s\in H^{G}_{k}(X)$,
$\frac{\lVert T^{G}_{k}[f]s\rVert^{2}}{\lVert s\rVert^{2}}$ $=\frac{\left(T^{G}_{k}[f]s\,\bigg{|}T^{G}_{k}[f]\,s\right)}{\int_{X}\langle s,\,s\rangle\,\mathrm{dV}_{X}}$ (16)
$=\frac{\lVert f\rVert^{2}_{\infty}\,\left(S^{G}_{k}[f]s\,\bigg{|}S^{G}_{k}[f]\,s\right)}{\int_{X}\langle s,\,s\rangle\,\mathrm{dV}_{X}}$
$\leq\lVert f\rVert^{2}_{\infty}\,.$
Hence, it proves the second inequality (15).
Now, we are ready to prove the theorem. Since $X$ is compact, then $\mu^{-1}(0)$ is closed and hence compact. Choose $x_{e}\in\mu^{-1}(0)$ a point with
$\lvert\widetilde{f}(x_{e})\rvert=\lVert f\rVert_{\infty}\,.$
Since the Berezin transform has as a leading term the identity, we have
$\left\lvert(I^{G}_{k}[f])(x_{e})-\widetilde{f}(x_{e})\right\rvert\leq\frac{C}{k}$
for a suitable $C>0$. Thus,
$\lVert f\rVert_{\infty}-\frac{C}{k}=\left\lvert\widetilde{f}(x_{e})\right\rvert-\frac{C}{k}\leq\left\lvert\left(I^{G}_{k}[f]\right)(x_{e})\right\rvert\leq\lVert I^{G}_{k}[f]\rVert_{\infty}\,.$
Putting (15) in this last inequality, we get part (a) of Theorem 1.1.
Now, part (b) and part (c) of Theorem 3.1 are consequence of the Theorem 2.3.
6 Proof of Corollary 1.1
Let $f$ and $g$ be two $G$ and $S^{1}$ invariant smooth function on $X_{G}$. For ease of notation we write $T^{G}_{k}[f]$ for $T^{G}_{k}[\widetilde{f}]$. It is clear by Theorem 1.1 and [GH1] that $\star$ is a well-defined associative star product. In fact, by part (b) of Theorem 1.1, we get
$\left\lVert T^{G}_{k}[{\{f,\,g\}-i(C_{1}(f,g)-C_{1}(g,f))}]\right\rVert=O\left(\frac{1}{k}\right)$
and taking the limit for $k\rightarrow+\infty$ and using part (a) we get
$\lVert\{f,\,g\}-i(C_{1}(f,g)-C_{1}(g,f))\rVert_{\infty}=0$
and thus we get part (b). Similarly, by Theorem 1.1 part (c) and part (a), we get that $C_{0}(g,h)=g\cdot h$. In a similar way the associativity property is proved.
Now, we prove the uniqueness. Let $C_{j}$ and $\tilde{C}_{j}$ be two bi-differential operators inducing a star product. Note $C_{0}=\tilde{C}_{0}$ by hypothesis. By induction assume that $C_{j}=\tilde{C}_{j}$ for each $j\leq N-2$. By the results in [GH1], we have
$T_{k}^{G}[f]\circ T_{k}^{G}[g]\sim\sum_{j=0}^{N-1}k^{-j}T_{k}^{G}[C_{j}(f,\,g)]+O(k^{-N})$
and
$T_{k}^{G}[f]\circ T_{k}^{G}[g]\sim\sum_{j=0}^{N-1}k^{-j}T_{k}^{G}[\tilde{C}_{j}(f,\,g)]+O(k^{-N})\,,$
where $\sim$ stand for “as the same asymptotic as”. For every $f,\,g\in C^{\infty}(X_{G})^{S^{1}}$, by the inductive hypothesis and subtracting these two expression, we obtain
$\left\lVert\frac{1}{k^{N-1}}\,T^{G}_{k}[{C}_{N-1}(f,\,g)-\tilde{C}_{N-1}(f,\,g)]\right\rVert\leq\frac{K}{k^{N}}\,.$
Eventually, taking the limit on both sides as $k$ goes to infinity and applying part a) of Theorem 1.1, we obtain
$C_{N-1}(f,\,g)=\tilde{C}_{N-1}(f,\,g)\,.$
The unit of the algebra $C^{\infty}(X_{G})^{S^{1}}$, the constant function $1$, will also be the unit in the star-product. In fact, it is equivalent to
[…]
for $j\geq 1$ and for every $f\in C^{\infty}(X_{G})^{S^{1}}$. First, by Theorem 2.1 note that
$C_{0}(1,\,g)=C_{0}(1,\,g)=g$
for every $k$. Now, recall that
$A_{N}=\hat{R}^{l}T^{G}[f]T^{G}[g]-\sum_{j=0}^{N-1}\hat{R}^{N-j}T^{G}[C_{j}(f,\,g)]$
which is an operator of Szegő type of order zero by Theorem 2.2 and it is invariant under the circle action. Thus, its symbol is a $G$ and $S^{1}$ invariant function on $X_{G}$ which, by restriction on $\mu^{-1}(0)$, defines a $S^{1}$-invariant function on $X_{G}$. By definition, it is the next element $C_{N}(f,\,g)$ in the star-product. Now, if we put $g=1$
$A_{1}=\hat{R}T^{G}[f]T^{G}[g]-\hat{R}T^{G}[f\cdot g]=0$
and hence the symbol of $A_{1}$ vanishes. By induction the claim (17) follows.
Let us now prove that the Berezin-Toeplitz star product fulfills parity. Considering the formal parameter to be real, $\nu=\overline{\nu}$, this is equivalent to
$\overline{C_{k}(f,\,g)}=C_{k}(\overline{g},\,\overline{f})\,,$ (18)
for $k\geq 0$. Note that, as an easy consequence of the definition of Berezin-Toeplitz star product, we have
$T_{k}^{G}[f]^{\dagger}=T_{k}^{G}[\overline{f}]\,.$
The star product $\overline{g}\star\overline{f}$ is given by the asymptotic expansion of the composition
$T_{k}^{G}\left[\overline{g}\right]\cdot T_{k}^{G}\left[\overline{f}\right]$ $=T_{k}^{G}[g]^{{\dagger}}\cdot T_{k}^{G}[f]^{{\dagger}}$
$=\left(T_{k}^{G}[{f}]\cdot T_{k}^{G}[g]\right)^{{\dagger}}$
$\sim\sum_{j=0}^{+\infty}k^{-j}T_{k}^{G}\left[C_{j}(f,\,g)\right]^{{\dagger}}$
$\sim\sum_{j=0}^{+\infty}k^{-j}T_{k}^{G}\left[\overline{C_{j}(f,\,g)}\right]\,.$
Thus, the claim (18) follows.
By [GH1], see also [GH2], the trace $\mathrm{Tr}_{k}$ on $H(X)^{G}_{k}$ admits a complete asymptotic expansion in decreasing integer power of $k$:
$\mathrm{Tr}_{k}\left(T^{G}_{k}[f]\right)\sim k^{d}\left(\sum_{k=0}^{+\infty}k^{-j}\,\tau_{j}(f)\right)\,,$
with $\tau_{j}(f)\in\mathbb{C}$. It induces a $\mathbb{C}[[\nu]]$-linear map
$\mathrm{Tr}\,:\,{C}^{\infty}(X_{G})^{S^{1}}[[\nu]]\rightarrow\nu^{-n}\,\mathbb{C}[[\nu]]$
such that
$\mathrm{Tr}(f):=\,\sum_{j=0}^{+\infty}\nu^{j-n}\,\tau_{j}(f)\,,$
where for $f\in C^{\infty}(M)$ the $\tau_{j}(f)$ are given by the asymptotic expansion above and for arbitrary elements by $\mathbb{C}[[\nu]]$-linear extension, $\nu=k^{-1}$. Now, we shall prove
$\mathrm{Tr}(f\star g)=\mathrm{Tr}(g\star f)\,.$ (19)
By $\mathbb{C}[[\nu]]$-linearity, we prove (19) for $f$ and $g$ in ${C}^{\infty}(X_{G})^{S^{1}}$. Note that $f\star g-g\star f$ is given by the asymptotic expansion of
$T^{G}_{k}[f]\,T^{G}_{k}[g]-T^{G}_{k}[g]\,T^{G}_{k}[f]\,.$
Hence the trace of $f\star g-g\star f$ is given by the expansion of
$\mathrm{Tr}_{k}\left(T^{G}_{k}[f]\,T^{G}_{k}[g]-T^{G}_{k}[g]\,T^{G}_{k}[f]\right)\,.$
But for every k this vanishes and thus we get (19).
Acknowledgments: The author thanks Chin-Yu Hsiao and Herrmann Hendrik for their valuable conversations. The author is a member of GNSAGA, part of the Istituto Nazionale di Alta Matematica, and expresses gratitude to the group for its support.